# Thinking Racial Bias in Fair Forgery Detection: Models, Datasets and Evaluations

### v1

@@ Abstract
Due to the successful development of deep image generation technology, forgery detection plays a more important role in social and economic security. Racial bias has not been explored thoroughly in the deep forgery detection field. In the paper, we first contribute a dedicated dataset called the Fair Forgery Detection (FairFD) dataset, where we prove the racial bias of public state-of-the-art (SOTA) methods. Different from existing forgery detection datasets, the self-construct FairFD dataset contains a balanced racial ratio and diverse forgery generation images with the largest-scale subjects. Additionally, we identify the problems with naive fairness metrics when benchmarking forgery detection models. To comprehensively evaluate fairness, we design novel metrics including Approach Averaged Metric and Utility Regularized Metric, which can avoid deceptive results. Extensive experiments conducted with nine representative forgery detection models demonstrate the value of the proposed dataset and the reasonability of the designed fairness metrics. We also conduct more in-depth analyses to offer more insights to inspire researchers in the community.
Keywords Face Forgery Detection  $\cdot$ Racial Bias  $\cdot$ Fairness Evaluation
1 Introduction
@@
Face forgery refers to the creation of fake images or videos of a person’s face using conventional techniques or deep learning methods. These forgeries can be used to spread misinformation, commit fraud, or even blackmail people. There are numerous methods are proposed for detecting face forgery [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Although an increasing number of advanced face forgery detection technologies are being developed, the racial fairness of these detectors is consistently overlooked by researchers [15]. Detectors with severe racial bias can lead to significant social impact. For example, these detectors might disproportionately label faces from a particular racial group as fake, thereby indicating discrimination towards this particular racial group. Therefore, when a detector is ready for deployment, evaluating and analyzing its fairness is a crucial process. However, although there is extensive available research about the fairness in machine learning to draw upon [16, 17, 18, 19, 20], evaluating the fairness in face forgery detection systems remains difficult. This is due to several distinct differences between face forgery detection and other machine learning tasks. The current conditions for evaluating the fairness of face forgery detection systems are still insufficient due to the following reasons:
Existing face forgery detection datasets are not suitable for racial fairness evaluation. We provide the statistical data for current mainstream datasets in Tab 2. And we propose three reasons why these existing datasets are not suitable for racial fairness evaluation. 1.Limited Number of Subjects. Current datasets usually have a limited number of subjects. However, we demonstrate that there is significant fluctuation in the model’s performance across subjects. This result implies these datasets are unsuitable for evaluating racial fairness since individual fairness may overshadow group fairness. Detailed analyses can be found in Sec 3.1. 2.Lack of Diversity in Forgery Approaches. We find that different forgery methods have different fairness levels (see Tab 1). So, we should strive to diversify forgery methods as much as possible, enabling a more comprehensive evaluation of the detectors’ fairness. Detailed description can be found in Sec 3.1. 3.Unbalanced Racial Ratio. Existing face forgery detection datasets are usually not race-balanced (see Tab 2).
@@
Xception Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
Caucasian Asian African Indian Mean STD AccGap
Real Face 0.7960 0.7201 0.8449 0.7715 0.7831 0.0450 0.1248
Fake Face FaceSwap 0.8910 0.8949 0.7668 0.8968 0.8624 0.0552 0.1300
SimSwap 0.3950 0.5436 0.2137 0.4264 0.3947 0.1183 0.3299
FaceReen 0.6102 0.6888 0.5223 0.6329 0.6136 0.0599 0.1665
DualReen 0.4114 0.5458 0.2406 0.4691 0.4167 0.1123 0.3052
MaskGAN 0.2593 0.3193 0.2013 0.2626 0.2606 0.0417 0.1180
Overall 0.5605 0.6188 0.4649 0.5766 0.5552 0.0563 0.1538
(a) Xception.
F3Net Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
Caucasian Asian African Indian Mean STD AccGap
Real Face 0.8117 0.7393 0.7958 0.7695 0.7791 0.0275 0.0724
Fake Face FaceSwap 0.9469 0.9454 0.8665 0.9510 0.9275 0.0352 0.0845
SimSwap 0.3391 0.4590 0.2032 0.3929 0.3486 0.0941 0.2558
FaceReen 0.5992 0.6775 0.6367 0.6773 0.6477 0.0325 0.0783
DualReen 0.1793 0.2838 0.1652 0.2036 0.2080 0.0459 0.1186
MaskGAN 0.2226 0.2643 0.2432 0.2127 0.2357 0.0198 0.0516
Overall 0.5165 0.5616 0.4851 0.5345 0.5244 0.0278 0.0765
(b) F3Net.
RECCE Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
Caucasian Asian African Indian Mean STD AccGap
Real Face 0.8423 0.7819 0.8799 0.8390 0.8358 0.0350 0.0980
Fake Face FaceSwap 0.8848 0.9036 0.7605 0.8867 0.8589 0.0573 0.1431
SimSwap 0.3054 0.4540 0.1547 0.3489 0.3158 0.1075 0.2993
FaceReen 0.5348 0.6161 0.4667 0.5632 0.5452 0.0539 0.1494
DualReen 0.2298 0.3533 0.1275 0.2583 0.2422 0.0805 0.2258
MaskGAN 0.2002 0.2552 0.1843 0.1947 0.2086 0.0275 0.0709
Overall 0.4996 0.5607 0.4289 0.5151 0.5011 0.0473 0.1318
(c) RECCE.
UCF Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
Caucasian Asian African Indian Mean STD AccGap
Real Face 0.8738 0.8008 0.8984 0.8752 0.8621 0.0367 0.0976
Fake Face FaceSwap 0.9192 0.9314 0.7722 0.9229 0.8864 0.0661 0.1592
SimSwap 0.3558 0.5081 0.1428 0.3488 0.3389 0.1299 0.3653
FaceReen 0.5889 0.7145 0.4831 0.6000 0.5966 0.0819 0.2314
DualReen 0.3756 0.5286 0.2420 0.4091 0.3888 0.1021 0.2866
MaskGAN 0.2615 0.3374 0.2130 0.2326 0.2611 0.0473 0.1244
Overall 0.5625 0.6368 0.4586 0.5648 0.5557 0.0635 0.1782
(d) UCF.
Table 1: Racial bias (STD and AccGap (the maximum differences in accuracy)) evaluations on four popular detectors (Xception [1], F3Net [12], RECCE [10], UCF [11]). The four detectors are trained on FF++ [1] and tested on our FairFD dataset. Training details can be found in Sec 5.1. The results are categorized by ethnicity (Caucasian, Asian, African, Indian). "Mean" donates the average accuracy for four ethnicities. The results are also categorized by real and fake faces. "Overall" represents the average accuracy of the total test dataset. Fake faces are split by forgery approaches.
Existing fairness metrics will cause deceptive results. We identify the bias offset and the aggregation distortion problem, which will cause deceptive racial bias. 1.Bias Offset Problem. Existing fairness metrics typically utilize overall average accuracy for calculations instead of separate calculations for each forgery method. However, this way may obscure part of the bias since different forgery methods may have different privileged races. A detailed explanation of the causes, examples, and solutions for the bias offset is in Sec 4.1. 2.Aggregation Distortion Problem. Current face forgery detectors often suffer from a significant limitation – they struggle to achieve satisfactory performance on unseen forgery methods. As a consequence, these models exhibit varying performance levels across different forgery techniques (See in Tab 1). For existing fairness metrics, even if two forgery methods yield the same bias, detectors with lower baselines are actually more unfair. Treating each forgery method as equal is unreasonable. Detailed explanation is in Sec 4.2.
This work aims to fill the gap in research on racial fairness in face forgery detection by proposing a comprehensive and accurate fairness evaluation system. To achieve this, we introduce new datasets and fairness metrics for assessing existing detectors. We present the workflow of fairness evaluation in Fig 1. The contributions of our work are summarized as follows:
@@
    To our knowledge, we first introduce the forgery detection racial bias evaluation benchmark which comprehensively provides the indicated large-scale dataset, fairness metrics and evaluation protocols.
@@
    We firstly introduce the Fair Forgery Detection (FairFD) dataset for racial bias evaluation, which contains the largest scale subjects, race-balanced ratio and incorporates diverse forgery approaches including FaceSwap [21], SimSwap [22], FastReen [23], DualReen [24], and MaskGAN [25].
@@
    We identify the bias offset and aggregation distortion problems with naive fairness metrics. Following, the novel Approach Averaged Metric and Utility Regularized Metric are designed to address the mentioned issues.
  • $\bullet$
    Extensive experimental results prove the limited fairness of SOTA methods and verify the value of the proposed metric. Besides, we also give in-depth analyses and interesting insights to promote the development of the field.
Dataset Race Rate Race Balanced Undefined Ethnicity Subject Number Approaches Real Frame Number Fake Frame Number
@@
ForgeryNet [30] - - - - - Yes 5400 15 1438.201k 1457.861k
FairFD(ours) $\sim$25$\sim$25$\sim$25$\sim$250 No 11430 5 52.041k 253.794k
@@
Table 2: Face Forgery Detection Dataset Comparison. Our dataset is race-balanced, with no undefined races, the number of subjects is maximized, and forgery approach exhibits diversity.
2 Related Work
2.1 Fairness in Face Forgery Detection
The research on fairness in face forgery detection is still relatively limited and waits for further exploration. A preliminary study [31] investigates bias in three commonly used face forgery detectors. They created real face images by sampling from the RFW [32] and UTKFace [33]. Next, they generate fake faces by blending two faces. However, they do not explicitly consider any fairness metrics. This simple study lacks a reasonable evaluation system but still verifies face forgery detectors have a significant racial bias to some extent. Study in [34] annotates five popular deepfake detection datasets with age, gender, ethnicity, etc. Due to the racial imbalance of current datasets, they also propose a metric to deal with the unbalanced test dataset. Another study [35] creat a gender-balanced dataset (GBDF) sampled from the FF++, Celeb-DF, and DF-1.0. This approach involves a limited number of subjects, which restricts the dataset’s capability for fairness evaluations.
Existing research on fairness in face forgery detection is evidently lacking depth. Although a gender-balanced dataset is proposed, it is derived from existing datasets, and our work reveals inherent limitations when using these datasets for evaluation. There is currently no suitable dataset to evaluate the fairness of face forgery detection systems. These limitations are overlooked in previous work. We are the first to conduct in-depth analyses of fairness in face forgery detection and develop a comprehensive and accurate fairness evaluation system.
2.2 Deepfake Datasets and Fairness Metrics
@@
Deepfake Dataset. We summarize the information of existing datasets in Tab 2. We provide the proportions of each race, along with whether the datasets are race-balanced. Additionally, we present whether undefined ethnicity faces are included(,i.e., faces from one race are replaced with another race. We also supply the number of subjects, the number of forgery approaches, and the total number of frames. We also give detailed description of these widely used face forgery datasets in Sec A.3 of Supp.
Fairness Metric. To evaluate racial fairness, what we require is a group fairness metric. There are various group fairness metrics, and the selection of a specific metric also depends on the application context. Commonly, researchers use the performance metric difference between privileged groups and unprivileged groups as a fairness metric. Demographic Parity Difference (DPD) [16, 17] utilizes the difference in positive rate, which represents the proportion of data predicted to be positive, as its fairness metric. The difference in Equalized Odds (DEOdds) [16] utilizes the average of the differences in true positive rate and false positive rate as its fairness metric. The difference in Equal Opportunity (DEO) [18] utilizes the difference in true positive rate solely as a fairness metric. In face recognition scenarios, the Standard Deviation (STD) of performance metrics across different groups is often used [32, 36, 37, 38, 39]. Equity-Scaled Segmentation Performance (ESSP) proposes to evaluate segmentation performance and group fairness simultaneously in medical image segmentation scenarios [19]. In the context of the generative model, the frequency of each group in the generated images is computed, and the average difference in frequency between each pair of groups is calculated as fairness metric [20]. Moreover, there are numerous works proposing more suitable metrics based on the application scenarios. In this work, we consider the commonly used metrics including DPD [16, 17], DEOdds [16], DEO [18], STD [32, 36, 37, 38, 39] and our proposed novel fairness metrics. The definitions of these metrics can be found in Sec A.2 of Supp.
3 FairFD Dataset
Figure 2: Representative examples. We select two images for each ethnicity. The top row represents the original image. Below this, from top to bottom, the forgery methods are FaceSwap [21], SimSwap [22], FastReen [23], DualReen [24], and MaskGAN [25].
3.1 Limitations of Current Datasets for Fairness Evaluation
The limitations of existing datasets are as follows:
@@ 3.1 Limitations of Current Datasets for Fairness Evaluation
Limited Number of Subjects. The construction process of the existing face forgery detection datasets involves collecting videos, subsequently creating forgeries, and then segmenting them into frames. As a result, there are typically a small number of subjects and each subject has a large number of frames in these datasets. But in fact, the limited number of subjects makes it challenging to draw meaningful comparisons across groups. Here we do a verification experiment to explain and prove it in Sec A.6 of Supp.
@@
Lack of Diversity in Forgery Approaches. In Tab 2, only DFDC and ForgeryNet employ a variety of forgery techniques. In our research, we find that different forgery methods have different fairness levels. We validate this point in the subsequent Tab 1. So, we should strive to diversify forgery methods as much as possible, enabling a more comprehensive evaluation of the system’s fairness.
@@
Unbalanced Racial Ratio. As stated in [32], a race-balanced test dataset helps ensure a fair evaluation across different race groups. An imbalanced test dataset may result in an unfair evaluation, where the performance of groups with a larger number can exert a more significant impact on fairness evaluation. In Tab 2, only DF-1.0 [29] has a balanced number of ethnic groups. Additionally, in identity-replaced forgery approaches, there is a possibility of faces from one ethnicity being replaced with those from another. In [34], this phenomenon is referred to as "undefined attribute annotation." Undefined attributes can also significantly lower the quality of the evaluation of racial fairness.
3.2 FairFD Description
@@
Having such limitations in existing datasets, we now introduce our dataset, FairFD, aiming to address these shortcomings. FairFD endeavors to overcome previous challenges and provide a more reasonable and comprehensive benchmark for evaluating fairness in face forgery detection. Representative examples of ours are presented in Fig 2. The overview of FairFD can be seen in the Tab 2. Subsequently, we delve into several pivotal facets of our dataset.
@@
Our dataset is an image-level dataset, consisting of original pristine images and 5 kinds of forgery images, i.e., FaceSwap [21], SimSwap [22], FastReen [23], DualReen [24], and MaskGAN [25]. Due to face forgery requires using detectors like Dlib [40] to identify faces and some faces are challenging to detect, the number of images of each forgery method is less than the number of pristine images. Specifically, the count of pristine images and the count of images for each forgery method are as follows: Pristine: 52.0k, FaceSwap: 50.9k, SimSwap: 46.7k, FastReen: 52.0k, DualReen: 52.0k, and MaskGAN: 52.0k. In addition to the forgery approach label, our approach also includes labels for four ethnicities (i.e., Caucasian, Asian, African, and Indian). Each ethnicity contains approximately 3000 subjects.
3.3 Source Data Collection
@@
To align with our requirements, which include having racial labels, achieving racial balance, and containing a sufficient number of subjects, we use the Racial Faces in the Wild(RFW) [32] dataset in face recognition filed as pristine images. The RFW dataset comprises face images with four racial labels (,i.e., Caucasian, Asian, African, and Indian), containing approximately 3000 subjects in each racial group, with a roughly equal distribution. Each subject has approximately $3\sim 7$ images. All images in the RFW dataset have a resolution of $400\times 400$ pixels. Besides, the images are carefully selected to maintain similar distributions in terms of age, gender, yaw angle, and pitch angle. RFW meets all our requirements. To reduce the time and human resources required for dataset collection and preprocessing, we directly use RFW as the source data.
3.4 Forgery Approach
To achieve the goal of diversity, we choose various approaches and techniques. We classify the forgery methods into face swap, expression reenactment and attribute manipulation. See details in Sec A.1 of Supp. 1) For face swap, we select FaceSwap [21] and SimSwap [22]. Although both methods result in a face-swapping effect, to ensure approach diversity, we select two distinct face-swapping approaches. FaceSwap [21] is a graphic-based face swap method. Whereas SimSwap [22] is a learning-based face swap method. 2) For expression reenactment, we select FastReen [23] and DualReen [24]. Both face swap and expression reenactment methods involve the use of source and target images to transfer faces or expressions from the target image to the source image. For each subject, we first designate it as the source subject. Next, within the same ethnic group, we randomly select another subject as the target subject, ensuring that the transfer occurs only within a single ethnic group as stated in Sec 3.1. Once the source-target pairs are established, these pairs remain fixed in other identity-replaced forgery approaches. Considering that each subject may have multiple images, for each specific image, we randomly select one image from its appointed target subject for the transfer, ensuring diversity and variability in the dataset. 3) For attribute manipulation, we select MaskGAN [25]. MaskGAN [25] provides an official GUI program that allows manual manipulation of attributes by making use of face parsing. However, as we aim to automate the face forgery process, we randomly select a subset from the nose, glasses, left eye, right eye, left eyebrow, right eyebrow, left ear, right ear, mouth, upper lip, and lower lip. We then apply random dilate and erode operations to the selected parts, enlarging or reducing the chosen regions. Missing parts are filled with skin, resulting in the final manipulated outcome. In summary, we select three types, a total of five forgery approaches, which achieve the diversity requirement.
4 The Proposed Evaluation Metrics
@@
Even though we have obtained a suitable dataset for evaluating deepfake detection’s fairness, we still can not get a credible evaluation result due to existing fairness metrics having two disadvantages. Firstly, the bias offset may lead to an underestimation of racial bias. Secondly, aggregation distortion can result in biased evaluation outcomes favoring specific forgery approaches. Below, we introduce the two disadvantages and their corresponding solutions respectively.
@@
4.1 Bias Offset Problem
@@
Bias offset refers to bias that will be partially obscured due to the calculation process of existing fairness metrics. Existing fairness metrics do not calculate separately for each forgery method instead of the final averaged performance scores. However, this way may obscure certain biases, which we call bias offset. Taking an example, in Fig 3, face forgery detectors may exhibit different biases for various forgery approaches. We calculate AccGap (the maximum differences in accuracy) and STD (standard deviation) for each forgery approach. In this example, both Forgery Approach 1 (FA1) and Forgery Approach 2 (FA2) exhibit an AccGap greater than 0.2 and an STD greater than 0.07. However, for Forgery Approach 1(FA1), the performance of Caucasians is better than Asians, while for FA2, the performance of Asians is better than Caucasians. In this situation, when calculating the fairness score using the final averaged performance scores, bias is to some extent offset, resulting in a smaller bias score. AccGap is less than 0.2, and the STD is less than 0.07 calculated using average accuracy. We refer to this phenomenon as Bias Offset. In Tab 1, there are some practical examples. Such as in UCF, for forgery approaches DualReen and MaskGAN, the accuracies of Caucasian and Indian will cause a Bias Offset problem. A more reasonable way is to calculate fairness metrics separately for each forgery method and average them. We call this fixed version as Approach Averaged Metric.
@@
$\text{AADPD}=\frac{1}{|\mathbb{F}|}\sum_{f\in\mathbb{F}}{\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}\vphantom{i}}}\left|P\left(\hat{Y}\,|\,S=s,F=f\right)-P\left(\hat{Y}\,|\,S=s^{\prime},F=f\right)\right|,$ (2)

### v3

@@ Abstract
Due to the successful development of deep image generation technology, forgery detection plays a more important role in social and economic security. Racial bias has not been explored thoroughly in the deep forgery detection field. In the paper, we first contribute a dedicated dataset called the Fair Forgery Detection (FairFD) dataset, where we prove the racial bias of public state-of-the-art (SOTA) methods. Different from existing forgery detection datasets, the self-constructed FairFD dataset contains a balanced racial ratio and diverse forgery generation images with the largest-scale subjects. Additionally, we identify the problems with naive fairness metrics when benchmarking forgery detection models. To comprehensively evaluate fairness, we design novel metrics including Approach Averaged Metric and Utility Regularized Metric, which can avoid deceptive results. We also present an effective and robust post-processing technique, Bias Pruning with Fair Activations (BPFA), which improves fairness without requiring retraining or weight updates. Extensive experiments conducted with 12 representative forgery detection models demonstrate the value of the proposed dataset and the reasonability of the designed fairness metrics. By applying the BPFA to the existing fairest detector, we achieve a new SOTA. Furthermore, we conduct more in-depth analyses to offer more insights to inspire researchers in the community. The source code is available at https://github.com/liudan193/Fairness-Benchmark-for-Face-Forgery-Detection.
Introduction
@@
Face forgery refers to the creation of fake images or videos of a person’s face using conventional techniques or deep learning methods. These forgeries can be used to spread misinformation, commit fraud, or even blackmail people. There are numerous methods are proposed for detecting face forgery [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Although an increasing number of advanced face forgery detection technologies are being developed, the racial fairness of these detectors is consistently overlooked by researchers [15]. Detectors with severe racial bias can lead to significant social impact. These detectors might disproportionately label faces from a particular racial group as fake, thereby indicating discrimination towards this particular racial group. Therefore, when a detector is ready for deployment, evaluating and analyzing its fairness is a crucial process. However, although there is extensive available research about the fairness in machine learning to draw upon [16, 17, 18, 19, 20], evaluating the fairness in face forgery detection systems remains difficult. This is due to several distinct differences between face forgery detection and other deep learning tasks.
This work aims to fill the gap in research on racial fairness in face forgery detection by proposing an accurate, comprehensive and credible fairness evaluation system. To achieve this goal, we analyze the shortcomings of existing evaluation components (i.e. dataset and metric), and our corresponding solutions. (1) Dataset. Existing face forgery detection datasets have a limited number of subjects. We find performance fluctuations significantly across subjects, so individual fairness may overshadow group fairness, which will make the evaluation results inaccurate. It also can be found that different forgery approaches have different fairness levels, limited forgery approaches will lead to a non-comprehensive result. Otherwise, undefined ethnicity (faces from two ethnicities are swapped) will lead to an inaccurate result. (2) Metric. We also propose two issues (Bias Offset and Aggregation Distortion) that will cause deceptive results. Bias Offset arises because existing fairness metrics typically use overall average accuracy for calculations instead of assessing each forgery method separately. This way may obscure some biases as different forgery methods may have different privileged races. Aggregation Distortion arises because detectors often show significant utility variations across different forgery techniques. Even if two forgery methods exhibit the same bias, detectors with lower utility can be more unfair. Treating each forgery method as equal will lead to unreliable results.
@@
To tackle these problems, we firstly introduce the Fair Forgery Detection (FairFD) dataset for racial bias evaluation, which contains the largest scale subjects, and incorporates diverse forgery approaches including Face Swapping: FaceSwap [21], SimSwap [22], Expression Reenactment: FastReen [23], DualReen [24], Face Editing: MaskGAN [25], StarGAN [26], StyleGAN [27], Diffusion-Based: SDSwap [28], DCFace [29], Face2Diffusion [30] and Transformer-Based: FSRT [31]. And the self-constructed FairFD dataset does not have any undefined ethnicity annotations. For the specific metric, we address the mentioned two issues by introducing the Approach Averaged Metric, which calculates fairness separately for each forgery approach and then aggregates them, and the Utility Regularized Metric, which uses the utility to regularize the fairness.
In addition to our evaluation system, we also introduce a new pruning approach called BPFA (Bias Pruning with Fair Activations). The designed BPFA leverages an innovative pruning metric to identify weights with the least impact on model utility while contributing most significantly to bias (e.g., racial bias). By pruning these weights, BPFA successfully enhances fairness without compromising utility. As a post-processing method, BPFA enhances fairness without any retraining. And it can be applied to any detector, including those already trained with existing fairness learning strategies, to further improve fairness performance. The workflow of fairness evaluation is illustrated in Figure 1. Sufficient experimental results prove that the proposed BPFA is an efficient, plug-and-play and robust pruning scheme, outperforming other baseline pruning methods by a significant margin.
The key contributions are summarized as follows:
@@
    To our knowledge, it is the early exploration to introduce a comprehensive racial bias evaluation benchmark for forgery detection, providing a large-scale dataset, fairness metrics, and evaluation protocols. We newly introduce the Fair Forgery Detection (FairFD) dataset for racial bias in forgery detection evaluation, which contains the largest scale subjects, race-balanced ratio and incorporates diverse forgery approaches.
@@
    We identify the bias offset and aggregation distortion problems with naive fairness metrics. Following, the novel Approach Averaged Metric and Utility Regularized Metric are designed to address the mentioned issues. Extensive experimental results demonstrate the limited fairness of existing SOTA methods and validate the value of our proposed metric.
@@
    We propose the Bias Pruning with Fair Activations (BPFA) method to improve the fairness of forgery detectors. Extensive experiments demonstrate the advantages of our method, particularly its efficiency, plug-and-play nature and robustness. By combining BPFA with the fairest detector, we achieve a new SOTA in racial fairness performance. Specifically, we offer in-depth analyses and insightful observations to advance the community.
Dataset Race Rate Race Balance Undefined Ethnicity Subject Number App- roach Real Img Number Fake Img Number
@@
ForgeryNet [36] - - - - - Yes 5400 15 1438k 1457k
FairFD(ours) $\sim$25$\sim$25$\sim$25$\sim$250 No 11430 11 52k 572k
@@
Table 1: Face Forgery Detection Dataset Comparison. Our dataset is race-balanced, with no undefined races, the maximized number of subjects and forgery approaches exhibit diversity.
Related Work
Fairness in Face Forgery Detection
Fairness Algorithm. Fairness in face forgery detection is a relatively novel topic. DAG-FDD [37] is first proposed to address fairness without demographic information by setting a probability threshold for minority groups to ensure low error rates for all groups meeting this threshold. DAW-FDD [37] utilizes demographic information to design losses to ensure similar performance across specified groups. PFGDFD [38] improves fairness by using disentanglement loss to separate demographic and forgery features.
@@
Deepfake Dataset. We summarize the information of existing datasets in Table 1. We provide the proportions of each race, along with whether the datasets are race-balanced. Additionally, we present whether undefined ethnicity faces are included(,i.e., faces from one race are replaced with another race. We also supply the number of subjects, the number of forgery approaches, and the total number of frames. DAG(W)-FDD [37] and PFGDFD [38] directly use several of the datasets mentioned above as test data. Our work reveals inherent limitations when using these datasets for evaluation. There is currently no suitable dataset to evaluate the fairness of forgery detection. We also give details of widely used face forgery datasets in the section "Face Forgery Detection Datasets" in Supp.
Fairness Metric. To evaluate racial fairness, what we require is group fairness metrics. There are various group fairness metrics, and the selection of a metric depends on the application context. In this work, we consider the commonly used metrics including DPD [16, 17], DEOdds [16], DEO [18], STD [39, 40, 41, 42, 43] and our proposed novel fairness metrics. The definitions of these metrics can be found in the section "Existing Fairness Metrics" in Supp.
FairFD Dataset
Limitations of Current Datasets
@@
Limited Number of Subjects. The construction process of the existing face forgery detection datasets involves collecting videos, subsequently creating forgeries, and then extracting frames. As a result, there are typically a small number of subjects and each subject has a large number of frames in these datasets. The limited number of subjects makes it challenging to draw meaningful comparisons across groups. Furthermore, we conducted the verification experiment to analyze and prove it in the section "Number of Subjects is Limited" in Supp.
@@
Lack of Diversity of Forgery Approaches. In Table 1, only DFDC and ForgeryNet employ a variety of forgery techniques. However, we find that different forgery methods have different fairness levels. We validate this point in the subsequent Figure 4. Thus, we should strive to diversify forgery methods as much as possible, enabling a more comprehensive evaluation of the system’s fairness.
@@
Undefined Attribute Annotation. For these identity-replaced forgery approaches, there is a possibility of faces from one ethnicity being replaced with those from another. In related work  [44], this phenomenon is referred to as "undefined attribute annotation." Undefined attributes can also significantly lower the quality of the evaluation of racial fairness, which is ignored in existing face forgery detection datasets.
FairFD Description
@@
Considering the mentioned limitations in existing datasets, we introduce our dataset, FairFD, aiming to address these shortcomings. FairFD endeavors to overcome previous challenges and provide a more accurate, reliable and comprehensive benchmark for evaluating fairness in face forgery detection. Representative examples of ours are presented in Figure 1. The overview of FairFD can be shown in the Table 1. Subsequently, we delve into several pivotal facets of our dataset.
@@
Our dataset is an image-level dataset, and for each image, there are 11 kinds of corresponding forgery images, i.e., Face Swapping: FaceSwap [21], SimSwap [22], Expression Reenactment: FastReen [23], DualReen [24], Face Editing: StarGAN [26], StyleGAN [27], MaskGAN [25], Diffusion-Based: SDSwap [28], DCFace [29], Face2Diffusion [30] and Transformer-Based: FSRT [31]. In addition to the forgery approach label, our approach also includes labels for four ethnicities (i.e., Caucasian, Asian, African, and Indian). Each ethnicity contains approximately 3000 subjects.
Source Data Collection and Forgery Process
@@
To align with our requirements, which include having racial labels, and containing a sufficient number of subjects, we use the RFW [39] dataset as pristine images. The RFW dataset comprises face images with four racial labels (,i.e., Caucasian, Asian, African, and Indian), containing approximately 3000 subjects for each racial group, with a roughly equal distribution. Each subject has approximately $3\sim 7$ images. All images in the RFW dataset have a resolution of $400\times 400$ pixels. Besides, the images are carefully selected to maintain similar distributions in terms of age, gender, yaw angle, and pitch angle. See details in "Detailed Distribution of FairFD" in Supp. To reduce the human resources, we use RFW as the source data.
To achieve the goal of diversity, we choose various approaches and techniques. We classify the forgery methods into face swap, expression reenactment, attribute manipulation and advanced forgery methods(stable diffusion and transformer). See details in the section "Classification of Forgery Approaches" in Supp. We reimplement these methods and apply them to the source data. Details configuration and process of forgery crafting can be found in the section "Forgery Crafting Process" in Supp.
The Proposed Evaluation Metrics
@@
Even though we have obtained a reliable dataset for evaluating deepfake detection’s fairness, we still can not get a credible evaluation result due to existing fairness metrics having two flaws. Firstly, the bias offset may lead to an underestimation of racial bias. Secondly, aggregation distortion can result in biased evaluation outcomes favoring specific forgery approaches. Below, we introduce the two flaws and their corresponding solutions respectively.
@@
Bias Offset Problem
@@
Bias offset refers to bias that will be partially obscured due to the calculation process of existing fairness metrics. Existing fairness metrics do not calculate separately for each forgery method instead of the final averaged performance scores. However, this way may obscure certain biases, which we call bias offset. Taking an example, in Figure 2, face forgery detectors may exhibit different biases for various forgery approaches. We calculate AccGap (the maximum differences in accuracy) and STD (standard deviation) for each forgery approach. In this example, both Forgery Approach 1 (FA1) and Forgery Approach 2 (FA2) exhibit an AccGap greater than 0.2 and an STD greater than 0.07. However, for Forgery Approach 1(FA1), the performance of Caucasians is better than Asians, while for FA2, the performance of Asians is better than Caucasians. In this situation, when calculating the fairness score using the final averaged performance scores, bias is to some extent offset, resulting in a smaller bias score. AccGap is less than 0.2, and the STD is less than 0.07 calculated using average accuracy. We refer to this phenomenon as bias offset. A more reliable way is to calculate fairness metrics separately for each forgery method and average them. We call this novel strategy as Approach Averaged Metric.
@@
AADPD $=\frac{1}{|\mathbb{F}|}\sum_{f\in\mathbb{F}}{\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}\vphantom{i}}}\bigl{|}P\left(\hat{Y}\,|\,S=s,F=f\right)$
$-P\left(\hat{Y}\,|\,S=s^{\prime},F=f\right)\bigr{|},$ (2)